OpenStudy (anonymous):

Can you find a and b so that the function f(x)= -tanx if xpi/4 is differentiable at x=pi/4? (It's greater than or equal to for the second part)

OpenStudy (xapproachesinfinity):

well you just apply derivative definition from left and right since it is differentiable at the point of the problem pi/4 then the two derivatives have to be equal

OpenStudy (noelgreco):

You still need this?

OpenStudy (anonymous):

Yes I do. So I would plug in pi/4 and for both equations and set them equal to each other?

OpenStudy (noelgreco):

It requires calculus, as you probably know. What would be the slope of the tangent function at x = pi/4?

OpenStudy (anonymous):

It would be the limit of the slope of the secant line.. which would be \[\lim h \rightarrow0 f(x _{1}+h)-f(x _{1}) / h\] But to find the slope, would I find the derivative at x=pi/4? Really confused

OpenStudy (xapproachesinfinity):

you don't need all that just do right derivative of f(x) and left derivative since f is differentiable at the point pi/4 the two derivative (right and left side of pi/4) should equal each other then solve for a,b you will need one more piece of information here f(pi/4) to get a